OpenStudy (anonymous):

A cooler contains 8 cans of soda's, 5 cokes,2 orange and 1 cherry. Two cans are slected at random without replacement. Find the probality that at leats one can is a coke. The prob that one can is a coke is?

OpenStudy (dumbcow):

Look at probability that none are cokes can 1: 3/8 because 3 cans are not cokes can 2: 2/7 you already took a non-coke leaving 2 out of the 7 remaining cans P(none) = 3/8*2/7 = 6/56 = 3/28 so P(at least 1) = 1 - 3/28

OpenStudy (anonymous):

8C2 = 28 means there are 28 possible ways to select two cans. Out of those selections, 3 selections do not contain a coke = 25/28, which is also 5 squared (5 cokes, 2 chances) over 8C2.

OpenStudy (anonymous):

so 25/28?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

r u guys like math teachers or something? u guys know everything!